1) How many excess electrons are required for an object to carry a charge of 1pC (10^-12 C)?

2) You remove 2x10^10 electrons from a quarter and the same number from a penny. Do the coins end up with a net positive charge or a net negative charge? which coin ends up with the greater net charge? explain.
Answer #1

1. n = Q/e = (10-12 C)/(1.6×10-19 C) = 6.25×106

2. Coins end up with positive charge. Both will have same charge. Charge on each is given by

Q = n e = (2×1010) ×(1.6×10-19 C) = 3.2×10-9 C